Cody is twice as old as Evan and five years ago the sum of their age was 14. Find the age of each boy now.
Answer by aquacean(7)


a. "scratchpad" method
b. "orthodox" method (algebra stuff, you know the stuff you show to your teachers)

A.
2 * 5 = 10
14 + 10 = 24
24 / 3 = 8
8 = Evan's present age
24 - 8 = 16
16 = Cody's present age
I multipied 5(years ago) by 2 (the product of which is 10) because they are two persons, thus each year is always multiplied by how many persons involved.
Then I added the 10 to 14 to find the total sum of their ages at present which is 24. And since Cody is twice as old as Evan, we divide 24 into 3, 1/3 of 24 is 8 which is Evan's age and 2/3 of it is 16 which is Cody's age.
B. "Orthodox" Method
a = Evan's age
2a = Cody's age
Evan's age:
3a = 14 + 5(2)
a = 24/3
a = 8
Cody's age:
2a = 8(2)
2a = 16